OpenStudy (anonymous):

how to solve k^2+3k=108

OpenStudy (anonymous):

k^2+3k-108=0 k^2+12k-9k-108=0 k(k+12)-9(k+12)=0 (k-9)(k+12)=0 k=9 and k=-12

OpenStudy (anonymous):

k^2 +3k - 108 = 0 (k+12)(k-9)=0 k=9 and k=-12

OpenStudy (anonymous):

try factoring as \[(x-9)(x+12)=0\]

OpenStudy (anonymous):

When the coefficient of x^2 is 1, then you can look for 2 numbers that add up to the x-coefficient of x and multiply together to give the constant. Often works with "made up" problems.

OpenStudy (anonymous):

hey estudier r u on facebook

OpenStudy (anonymous):

Yes, just so I can log in other places without any headaches..

OpenStudy (anonymous):

lolx hey can u give ur id so that i can add u as a friend

OpenStudy (anonymous):

Like I said, I don't use the account, so it's a bit pointless...sorry :-)

OpenStudy (anonymous):

oh................np